An electron and a proton are 0.53 x 10^-10 m apart.  Calculate their dipole moment if they are at rest. Using your dipole moment calculation, find the electric potential at a point 1.1 x 10^-9 m away when this point is along the axis of the dipole nearer the positive charge and when this point is 45o left of the axis but nearer the negative charge.
Answer #1

If they are at rest:

The answer would be the magnitude of the charge times the distance between them.

Charge of electron -= 1.6*10^-19

Distance= 0.53*10^-10m

So at rest the moment would be: (1.6*10^-19)*(0.53*10^-10)= 8.48*10^-30 C*m